MATH 172 Section 502  Calculus II  P. Yasskin Fall 1995

INFORMATION SHEET

INSTRUCTOR:  Dr. Philip B. Yasskin
OFFICE:  Blocker 620 I
OFFICE HOURS: MW 10:15-11:15 in BLOC 620 I

Th 10:25-11:00 in BLOC 126 or by appointment

OFFICE PHONE: 845-3734
MESSAGES:  845-3261 LEAVE YOUR PHONE NUMBER!
E-MAIL:  yasskin@math.tamu.edu
LECTURES:  MW 1:50 – 2:40 BLOC 156
Th 2:20 – 3:10 ZACH 128B
COMPUTER LAB: F  1:50 – 2:40 BLOC 126
REQUIRED TEXTS: Stewart – Calculus, 3rd Edition

Boggess, Barrow, Belmonte, Massoud, Morgan, Rahe, Stecher, Yasskin – CalcLabs with Maple V

RECOMMENDED SOFTWARE:

Maple V, Release 3, Student Edition with Maple Flight Manual

DESCRIPTION: This is a second course in calculus for math and science majors and covers Chs. 6 through 10 and Secs. 15.1-15.2 of Stewart. In addition we will use a symbolic computer language called MAPLE.

GRADING:  POINTS:  DATES:
Exam 1  150  Th 10/3 7:30PM – 9:30PM
Exam 2  150  Tue 10/29 7:30PM – 9:30PM
Exam 3  200  Tue 12/3 7:30PM – 9:30PM
Maple Quiz 1  50  To Be Announced
Maple Quiz 2  50  To Be Announced
Maple Project  50  To Be Announced
Final  250  Tue 12/17, 3:30PM – 5:30PM, BLOC 156
HW, Quiz & Lab  100
TOTAL  1000

I may curve any grade or the total and will then compute the course grade from the following table:

A = 900 – 1000 points  D = 600 – 699 points
B = 800 – 899 points  F = 0 – 599 points
C = 700 – 799 points

CALALOGUE DESCRIPTION:

172. Calculus. (4-0). Credit 4. I, II Techniques of integration, sequences, indeterminant forms, improper integrals, transcendental functions, parametric equations, polar coordinates, infinite series. Prerequisite: MATH 151 or 171.

OTHER POLICIES

1. HOMEWORK will be collected at the assigned class period. Some problems may

be graded but you will not know which ones. Late homework will NOT be accepted.

QUIZZES will be given in class and may not be announced. There will be NO make-ups for quizzes. On some quizzes, you may work with a few other students as a team and your grade will depend on the performance of the whole team.

LAB REPORTS will be collected at the lab period one week after the lab. Late lab reports will NOT be accepted. On lab reports, students will work in pairs. Each pair will turn in one lab report and receive one grade.

Homeworks, quizzes and lab reports will each count equally. The lowest two homework or quiz or lab report grades will be dropped. The remaining grades will be averaged and then rescaled to 100 points.

2. MAKE-UPS for MAJOR EXAMS will be given only in case of an absence authorized under University Regulations. You will need a note from your doctor or your academic advisor. If you know in advance that you will miss an exam, please contact me so that you can take the make-up in advance. If you have trouble reaching me, leave a message with the Math department secretaries (845-3261) and be sure to leave your phone number.

3. ATTENDANCE is REQUIRED. Attendance will be taken in lecture. If you sign the roll sheet, you are expected to remain in the classroom for the entire 50 minutes. More than 2 absences may have a detrimental effect on your grade especially in borderline cases.

4. You will be asked to provide multiple choice SCANTRON forms. You must have your ID with you at all exams. You may use a CALCULATOR on half of each exam. You will use the COMPUTER during Maple quizzes.

5. All material handed out or written on the board or spoken in class is copyrighted by the instructor.